the sum of the number and its square root is 6/25,then find the number ?

: the sum of the number and its square root is 6/25,then find the number ?


Answer by shahid(44)

let require number be x then following the conditions we get an equation
x+sqrt(x)=6/25 now substitute sqrt(x)=y
then x =y^2 by squaring both sides
y^2+y=6/25
25y^2+25y=6 25 is multiplied on L.H.S
25y^2+25y-6=0
25y^2+30y-5y-6=0
5y(5y+6)-1(5y+6)=0
(5y+6)(5y-1)=0
either 5y+6=0 or 5y-1=0
5y=-6 5y=1
y=-6/5 y=1/5
but from our substitution x=y^2
either x=(-6/5)^2 or x=(1/5)^2
x=36/25 or x=1/25 hope this is helpful fo you.
Question 143695: the sum of the number and its square root is 6/25,then find the number ?

: the sum of the number and its square root is 6/25,then find the number ?


Answer by jim_thompson5910(8196)

The previous solution has one flaw. There is only one solution and the answer is 1/25


Check:
1/25+sqrt(1/25)=1/25+1/5=1/25+5/25=6/25